A rectangular prism has a surface area of 448 cm2. Its length is 14cm and its width id 6cm find its height
Answer by Cromlix(4381)

Hi there,
The surface area of a
rectangular prism with
Length = 14cm
Width = 6cm
Height = hcm
Considering the 6 faces.
2 x 6 x h (Ends)
2 x 6 x 14 (Bottom and top)
2 x 14 x h (Sides)
Totalled up:
12h + 168 + 28h = 448 cm^2
40h + 168 = 448 cm^2
40h = 448 - 168
40h = 280
h = 7cm
Height = 7cm
Hope this helps:-)